# A Multi-Modal Explainability Approach for Human-Aware Robots in Multi-Party Conversation

### v1

@@ Acknowledgments
Table B.2: Architecture of our final IAE model. All convolution layers use kernel size 5. Relevant layer-specific hyper-parameters are listed in the brackets.
@@
Table B.3: Hyperparameters of the XAE model. Parameters with indices 1 – 4 control Vit (M1), MLP for pose vectors, intermediate network (M2), and GRU with attention (M3) respectively.
XAE model hyperparameters
@@
Table B.4: Hyperparameters/architecture of our final XAE model. For processing face images, we use the standard vision transformer architecture, therefore, we only provide the chosen hyperparameters. The architecture of the recurrent part is discussed in more detail in Paragraph 3.1.2.

### v2

@@ 3.1.4 Training and performance evaluation
The number of hyper-parameters optimized for the XAE model was much lower. They were optimized in two groups. The first group contained optimization-related parameters, such as gammas and learning rates, the second group contained architecture-related parameters. The resulting values are in Tables 3 and 4, while the latter mentioned also contains information about the network architecture. Activation function in the recurrent network is used on inputs to the GRU unit and outputs from the GRU (i.e., inputs to the final fully-connected layer).
Table 2: Architecture of our final IAE model. All convolution layers use kernel size 5. Relevant layer-specific hyper-parameters are listed in the brackets.
@@
Table 3: Hyper-parameters of the XAE model. Parameters with indices 1 – 4 control Vit (M1), MLP for pose vectors, intermediate network (M2), and GRU with attention (M3) respectively.
XAE model hyper-parameters
@@
Table 4: Hyper-parameters/architecture of our final XAE model. For processing face images, we use the standard vision transformer architecture, therefore, for brevity, we only provide the chosen hyper-parameter values. The architecture of the recurrent part is discussed in more detail in Paragraph 3.1.2.